# Density-Aware Farthest Point Sampling

### v1

@@ 5.1 Density estimation
$\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}({\bm{x}}):=\frac{\sum\limits_{{\bm{x}}_{i}\in\mathcal{D}_{\mathcal{X}}}K\left(\frac{{\bm{x}}-{\bm{x}}_{i}}{r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})}\right)}{|\mathcal{D}_{\mathcal{X}}|\left(r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})\right)^{d}}\enspace\text{ and }\enspace\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}}):=\frac{\sum\limits_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}K\left(\frac{{\bm{x}}-{\bm{x}}_{j}}{r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})}\right)}{|\mathcal{L}_{\mathcal{X}}|\left(r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})\right)^{d}},$ (5)
@@
$r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}}):=\min\left\{\min_{\bar{{\bm{x}}}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-\bar{{\bm{x}}}\|_{2}+\frac{\epsilon_{\mathcal{X}}}{|\mathcal{L}_{\mathcal{X}}|},\rho_{k}({\bm{x}})\right\}\enspace\text{ and }\enspace K({\bm{x}}):=\begin{cases}\frac{1}{V_{d}},&\text{if }\|{\bm{x}}\|_{2}\leq 1,\\ 0,&\text{otherwise,}\end{cases}$ (6)
with $\rho_{k}({\bm{x}}):=\min\{r\in\mathbb{R}^{+}\text{ such that }|\{\tilde{{\bm{x}}}\in\mathcal{D}_{\mathcal{X}}\mid\|{\bm{x}}-\tilde{{\bm{x}}}\|_{2}\leq r\}|\geq k\}$ the distance between $\bm{x}$} and its $k$-nearest neighbor in $\mathcal{D}_{\mathcal{X}}$, and $V_{d}:=\frac{\pi^{d/2}}{\Gamma(d/2+1)}$ is the volume of the unit ball in $\mathbb{R}^{d}$, where $\Gamma(r)=(r-1)!$ is the so-called Gamma function. The quantity $r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})$ defines an adaptive neighborhood size, depending on the selected set $\mathcal{L}_{\mathcal{X}}$, aiming at reducing the estimation bias at finite sample sizes (Wang et al., 2009). $\epsilon_{\mathcal{X}}$ in the definition of $r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})$ is an arbitrary small positive scalar that prevents the denominator in the kkNN density estimations from becoming zero on points in $\mathcal{L}_{\mathcal{X}}$. Note that, we are interested in estimating the densities only for points in $\mathcal{D}_{\mathcal{X}}\backslash\mathcal{L}_{\mathcal{X}}:=\{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}\text{ such that }{\bm{x}}\notin\mathcal{L}_{\mathcal{X}}\}$, which are not already in $\mathcal{L}_{\mathcal{X}}$, and we may want to select depending on the value they would provide, which we quantify with the associated weighted distance. Nonetheless, the kernel estimations in (5) are well-defined for all points in $\mathcal{X}$. In Appendix D we show that, under some assumptions, $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}$ and $\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}$ are asymptotically unbiased estimations of $p_{\mathcal{X}_{\mathcal{L}}}$ and $p_{\mathcal{X}_{\mathcal{D}}}$, respectively.
@@ 5.2 Estimated weighted fill distance
To see this consider that the numerator of $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})$, as defined in (5), is either $\frac{1}{V_{d}}$ or zero depending on whether there exists an ${\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}\cap\mathcal{N}_{k}({\bm{x}})$ or not. Depending on the ratio $\frac{|\mathcal{D}_{\mathcal{X}}|}{\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})|\mathcal{L}_{\mathcal{X}}|}$ and the value of $k$, the estimated weight function in (9) may allow for negative values on data points in $\mathcal{D}_{\mathcal{X}}\backslash\mathcal{L}_{\mathcal{X}}$, not yet included in the training set, where we would expect $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\leq\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. This behavior is not consistent with the behavior of the true weight function defined in (3), which associates non-negative values with points where $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\leq p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. This artifact is directly linked to the fact that the training and data densities are estimated using sets with different amount of points, i.e., if $|\mathcal{D}_{\mathcal{X}}|=|\mathcal{L}_{\mathcal{X}}|$ the estimated weight function in (9) would be non-negative. Thus, the approach we use to compute the densities’ ratio is affected by scaling issues related to the sets’ imbalance. But, we observe that the values of the estimated weight function are directly correlated with those of the weights $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ defined in (8). Given $\tilde{{\bm{x}}},\bar{{\bm{x}}}\in\mathcal{X}$ we have $\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\tilde{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\tilde{{\bm{x}}})}\right)>\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\bar{{\bm{x}}})}\right)\Rightarrow\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\tilde{{\bm{x}}})\geq\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\bar{{\bm{x}}})$ and $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\tilde{{\bm{x}}})>\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\bar{{\bm{x}}})\Rightarrow\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\tilde{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\tilde{{\bm{x}}})}\right)>\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\bar{{\bm{x}}})}\right).$
@@
The value of $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ is simply the number of points in $\mathcal{D}_{\mathcal{X}}$ that are close to $\bm{x}$}. Based on these observations, we avoid the scaling issue of the estimated weight function in (9) by replacing it with weights defined in (8). That is, instead of attempting the minimization of the quantity in (7) we consider the following minimization problem:
@@ 6 Simple illustration of DA-FPS
We provide a simple example to illustrate how DA-FPS works compared to uniform random sampling and Farthest Point Sampling (FPS). Fig. 1 (top row) shows a dataset of 1000 two-dimensional points in the unit square. The dataset contains a high-density cluster (650 points) in the center, a smaller cluster (200 points) in the lower left, and 150 points scattered uniformly at random. The blue background is a 2D kernel density estimation (KDE) plot, where darker blue means higher density. For the KDE plot we use the Seaborn Python library (Waskom, 2021). The second row of Fig. 1 shows 100-point subsets selected by uniform random sampling, FPS, and DA-FPS. For DA-FPS, we set $u=0$ and $k=100$. From a global perspective, DA-FPS balances two objectives: covering the data space evenly while giving greater importance to higher-density regions. Consequently, the sampled data approximates the original distribution but with reduced density differences across the data space. In contrast, FPS consistently selects points to provide a uniform representation of the data space, while uniform random sampling mainly focuses on the high-density clusters, potentially neglecting sparser regions, e.g., no point is selected in the lower right corner when uniform random sampling is used. Locally, both FPS and DA-FPS tend to select points that are evenly distributed without clustering. This is due to their optimization processes, which lead to maximize pairwise (weighted) distances between the selected points, creating a “repulsion effect”. One of the key advantages of DA-FPS for sampling training sets lies in its ability to balance the representation of both dense and sparse regions in the data. By ensuring that high-density regions are well-represented while also covering sparser areas, DA-FPS mitigates the risk of over-fitting or under-fitting to dominant clusters, which is a common issue when using random sampling and FPS, respectively.
@@ 7 Analysis of DA-FPS
This section analyzes how well DA-FPS approximates the solution to the optimization problem in (10). The following theorem establishes that DA-FPS achieves a $2k$-approximation for this problem, where $k$ is the number of nearest neighbors used in DA-FPS.
@@
Proof is in appendix˜E. Theorem 7.1 gives an upper bound on the approximation error of DA-FPS for the estimated weighted fill distance minimization problem defined in (10). The theorem indicates that increasing the number $k$ increases the approximation error. Recall that, $k$ is the largest possible values of the weights $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$. The presence of $k$ in the bound quantifies the worst case approximation error of the weights values, which depend on the selected set $\mathcal{L}_{\mathcal{X}}$. In particular, the weights associated with the set selected by DA-FPS can differ significantly from those of an optimal set $\mathcal{O}_{\mathcal{X}}$ of size $b$, especially in the early stages of sampling when $\mathcal{L}_{\mathcal{X}}$ contains only a few points. Appendix N provides an additional result stating that the approximation error of the DA-FPS estimation depends on the relationship between the weights of an optimal set and those of the set selected by DA-FPS.
It is important to note that the bound in Theorem 7.1 does not account for the quality of the density ratio estimation itself, which also depends on the parameter $k$. In other words, it does not quantify how well the estimated weighted fill distance in (11) relates to the true weighted fill distance in (2). Addressing this issue is an open problem for future work and is closely related to the broader field of density ratio estimation. Nonetheless, multivariate density estimation theory indicates that smaller selected sets lead to more biased approximations of the corresponding density marginal in (5), especially for high-dimensional data (Wang et al., 2009). In simpler terms, the smaller the size of the selected set, the less representative are the computed weights of the actual distribution of the points. To mitigate these limitations, we make use of the hyperparameter $u>0$. During the initial steps of the sampling process we consider constant weights, which we then iteratively update according to (8) after selecting the first $u$ points. Thus, in early stages, DA-FPS coincides with FPS. That is, it focuses on minimizing the maximum distance between any point in the dataset $\mathcal{D}_{\mathcal{X}}$ and its closest selected element. This approach ensures broader coverage of the data space without initially accounting for the density of the data. The value of $u$ plays a crucial role in the effectiveness of DA-FPS for training data selection in regression task involving high-dimensional data. If $u$ is set too small, sparse regions of the dataset may be underrepresented in the training set, as the sampling will primarily target higher-density regions. This imbalance can result in low model performances in sparse regions of the dataset with a consequential increase in the average error.

### v2

@@ 5.1 Density estimation
$\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}({\bm{x}}):=\frac{\sum\limits_{{\bm{x}}_{i}\in\mathcal{D}_{\mathcal{X}}}K\left(\frac{{\bm{x}}-{\bm{x}}_{i}}{r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})}\right)}{|\mathcal{D}_{\mathcal{X}}|\left(r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})\right)^{d}}\kern 5.0pt\text{ and }\kern 5.0pt\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}}):=\frac{\sum\limits_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}K\left(\frac{{\bm{x}}-{\bm{x}}_{j}}{r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})}\right)}{|\mathcal{L}_{\mathcal{X}}|\left(r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})\right)^{d}},$ (5)
@@
$r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}}):=\min\left\{\min_{\bar{{\bm{x}}}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-\bar{{\bm{x}}}\|_{2}+\frac{\epsilon_{\mathcal{X}}}{|\mathcal{L}_{\mathcal{X}}|},\rho_{k}({\bm{x}})\right\}\kern 5.0pt\text{ and }\kern 5.0ptK({\bm{x}}):=\begin{cases}\frac{1}{V_{d}},&\text{if }\|{\bm{x}}\|_{2}\leq 1,\\ 0,&\text{otherwise,}\end{cases}$ (6)
with $\rho_{k}({\bm{x}}):=\min\{r\in\mathbb{R}^{+}\text{ such that }|\{\tilde{{\bm{x}}}\in\mathcal{D}_{\mathcal{X}}\mid\|{\bm{x}}-\tilde{{\bm{x}}}\|_{2}\leq r\}|\geq k\}$ the distance between $\bm{x}$} and its $k$-nearest neighbor in $\mathcal{D}_{\mathcal{X}}$, and $V_{d}:=\frac{\pi^{d/2}}{\Gamma(d/2+1)}$ is the volume of the unit ball in $\mathbb{R}^{d}$, where $\Gamma(r)=(r-1)!$ is the so-called Gamma function. The quantity $r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})$ defines an adaptive neighborhood size, depending on the selected set $\mathcal{L}_{\mathcal{X}}$, aiming at reducing the estimation bias at finite sample sizes (Wang et al., 2009). $\epsilon_{\mathcal{X}}$ in the definition of $r_{\mathcal{L}_{\mathcal{X}}}^{k}({\bm{x}})$ is an arbitrarily small positive scalar that prevents the denominator in the kkNN density estimations from becoming zero on points in $\mathcal{L}_{\mathcal{X}}$. Note that, we are interested in estimating the densities only for points in $\mathcal{D}_{\mathcal{X}}\backslash\mathcal{L}_{\mathcal{X}}:=\{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}\text{ such that }{\bm{x}}\notin\mathcal{L}_{\mathcal{X}}\}$, which are not already in $\mathcal{L}_{\mathcal{X}}$, and we may want to select depending on the value they would provide, which we quantify with the associated weighted distance. Nonetheless, the kernel estimations in (5) are well-defined for all points in $\mathcal{X}$. In Appendix D we show that, under some assumptions, $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}$ and $\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}$ are asymptotically unbiased estimations of $p_{\mathcal{X}_{\mathcal{L}}}$ and $p_{\mathcal{X}_{\mathcal{D}}}$, respectively.
@@ 5.2 Estimated weighted fill distance
To see this consider that the numerator of $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})$, as defined in (5), is either $\frac{1}{V_{d}}$ or zero depending on whether there exists an ${\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}\cap\mathcal{N}_{k}({\bm{x}})$ or not. Depending on the ratio $\frac{|\mathcal{D}_{\mathcal{X}}|}{\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})|\mathcal{L}_{\mathcal{X}}|}$ and the value of $k$, the estimated weight function in (9) may allow for negative values on data points in $\mathcal{D}_{\mathcal{X}}\backslash\mathcal{L}_{\mathcal{X}}$, not yet included in the training set, where we would expect $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\leq\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. This behavior is not consistent with the behavior of the true weight function defined in (3), which associates non-negative values with points where $p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\leq p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$. This artifact is directly linked to the fact that the training and data densities are estimated using sets with different amount of points, i.e., if $|\mathcal{D}_{\mathcal{X}}|=|\mathcal{L}_{\mathcal{X}}|$ the estimated weight function in (9) would be non-negative.
Thus, the approach we use to compute the densities’ ratio is affected by scaling issues related to the sets’ imbalance. But, we observe that the values of the estimated weight function are directly correlated with those of the weights $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ defined in (8). Given $\tilde{{\bm{x}}},\bar{{\bm{x}}}\in\mathcal{X}$ we have $\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\tilde{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\tilde{{\bm{x}}})}\right)>\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\bar{{\bm{x}}})}\right)\Rightarrow\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\tilde{{\bm{x}}})\geq\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\bar{{\bm{x}}})$ and $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\tilde{{\bm{x}}})>\omega^{k}_{\mathcal{L}_{\mathcal{X}}}(\bar{{\bm{x}}})\Rightarrow\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\tilde{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\tilde{{\bm{x}}})}\right)>\left(1-\frac{\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})}{\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}(\bar{{\bm{x}}})}\right).$
@@
The value of $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ is simply the number of points in $\mathcal{D}_{\mathcal{X}}$ that are close to $\bm{x}$}. Based on these observations, we avoid the scaling issue of the estimated weight function in (9) by replacing it with weights defined in (8).
That is, instead of attempting the minimization of the quantity in (7) we consider the following minimization problem:
@@ 5.3 DA-FPS algorithm
The behavior of the algorithm is primarily controlled by two hyperparameters: the neighborhood size $k$ for the kkNN density estimation and the parameter $u$ for initial uniform selection. The parameter $k$ controls how we estimate density and directly affects the weight values $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$. Small values of $k$ make the weights very responsive to local changes in density, which may lead to high variance in the values of the weights, and consequently, to noticeable differences in the values of weights associated with points lying in regions with similar true density. Large values of $k$ produce more stable weights but may over-smooth important density variations. As we select more points, the impact of the parameter $k$ becomes less relevant because the weight value at a given location is increasingly determined by how close the nearest selected point is, rather than by the distance from the $k$-th nearest neighbor. This happens because each weight value counts points within a neighborhood whose size depends on the distance to the closest selected point. The parameter $u$ determines the amount of points selected with uniform weight: setting $u=0$ recovers pure density-aware selection from the start, while $u=b$ reduces to standard FPS. In Section 8, where we report experimental results, we comment on the influence of $k$ and $u$ on DA-FPS performance.
Note that we are considering the Euclidean distance in our formulation of the (estimated) weighted fill distance. However, other distance metrics can be used depending on the data characteristics and the specific application. We follow common practice and use the Euclidean distance for general-purpose applications. We remark that alternative formulations for the estimated weighted fill distance could be considered during the development of DA-FPS. For instance, we chose multiplicative weighting ($\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}$) rather than additive weighting ($\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}+\alpha\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$) because multiplicative weighting was motivated by our theoretical result and avoids the need to tune an additional scaling parameter $\alpha$. Further investigation into alternative formulations and their empirical performance could yield valuable insights and potentially lead to improved sampling strategies.
@@ 6 Simple illustration of DA-FPS
In this section we provide an intuitive understanding of DA-FPS behavior through a visual 2D example. We compare DA-FPS with uniform random sampling and standard FPS on a synthetic dataset with varying density regions. The illustration demonstrates how DA-FPS balances coverage of the feature space with density awareness, selecting points from high-density regions while still covering sparse areas.
@@
Fig. 1 (top row) shows a dataset of 1000 two-dimensional points in the unit square. The dataset contains a high-density cluster (650 points) in the center, a smaller cluster (200 points) in the lower left, and 150 points scattered uniformly at random. The blue background is a 2D kernel density estimation (KDE) plot, where darker blue means higher density. For the KDE plot we use the Seaborn Python library (Waskom, 2021). The second row of Fig. 1 shows 100-point subsets selected by uniform random sampling, FPS, and DA-FPS. For DA-FPS, we set $u=0$ and $k=100$. From a global perspective, DA-FPS balances two objectives: covering the data space evenly while giving greater importance to higher-density regions. Consequently, the sampled data approximates the original distribution but with reduced density differences across the data space. In contrast, FPS consistently selects points to provide a uniform representation of the data space, while uniform random sampling mainly focuses on the high-density clusters, potentially neglecting sparser regions, e.g., no point is selected in the lower right corner when uniform random sampling is used. Locally, both FPS and DA-FPS tend to select points that are evenly distributed without clustering. This is due to their optimization processes, which lead to maximize pairwise (weighted) distances between the selected points, creating a “repulsion effect”. One of the key advantages of DA-FPS for sampling training sets lies in its ability to balance the representation of both dense and sparse regions in the data. By ensuring that high-density regions are well-represented while also covering sparser areas, DA-FPS mitigates the risk of over-fitting or under-fitting to dominant clusters, which is a common issue when using random sampling and FPS, respectively.
@@ 7 Analysis of DA-FPS
In this section we provide a theoretical analysis of DA-FPS performance. We establish that DA-FPS achieves a $2k$-approximation for the estimated weighted fill distance minimization problem in (10), where $k$ is the number of nearest neighbors used in DA-FPS. Following is a discussion of the relationship between the parameter $k$ and approximation quality, as well as computational complexity considerations. We also address limitations of our density estimation approach and implementation details. The following theorem establishes that DA-FPS achieves a $2k$-approximation for the fill distance minimization problem in (10).
@@
Proof is in Appendix E. Theorem 7.1 gives an upper bound on the approximation error of DA-FPS for the estimated weighted fill distance minimization problem defined in (10). The theorem indicates that increasing the number $k$ increases the approximation error. Recall that $k$ is the largest possible value of the weights $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$. The presence of $k$ in the bound quantifies the worst case approximation error of the weights values, which depend on the selected set $\mathcal{L}_{\mathcal{X}}$. In particular, the weights associated with the set selected by DA-FPS can differ significantly from those of an optimal set $\mathcal{O}_{\mathcal{X}}$ of size $b$, especially in the early stages of sampling when $\mathcal{L}_{\mathcal{X}}$ contains only a few points. Appendix N provides an additional result stating that the approximation error of the DA-FPS estimation depends on the relationship between the weights of an optimal set and those of the set selected by DA-FPS.
It is important to note that the bound in Theorem 7.1 does not account for the quality of the density ratio estimation itself, which also depends on the parameter $k$. In other words, it does not quantify how well the estimated weighted fill distance in (11) relates to the true weighted fill distance in (2). Addressing this issue is an open problem for future work and is closely related to the broader field of density ratio estimation.
The estimated weighted fill distance in (11) may not accurately reflect the true weighted fill distance in (2), particularly when the training set is small or the data are high-dimensional. According to multivariate density estimation theory (Wang et al., 2009), limited sample sizes lead to biased density estimates in (5), which in turn cause the behavior of the estimated weights $\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ to deviate from that of the true weights $\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ in (3). Consequently, DA-FPS may appear effective for the estimated problem in (10) but may fail to provide reliable performance for the true problem, especially for small training sets under high-dimensional settings where density estimation is challenging. To mitigate these limitations, we first sample attempting to ensure broad coverage of the data space by using the hyperparameter $u>0$. That is, during the initial steps of the sampling process, when density estimates are unreliable, we consider constant weights, which we then iteratively update according to (8) after selecting the first $u$ points. Thus, in early stages, DA-FPS coincides with FPS and focuses on minimizing the maximum distance between any point in the dataset $\mathcal{D}_{\mathcal{X}}$ and its closest selected element. We note that, the theoretical approximation guarantee in Theorem 7.1 applies only to $u=0$. Nonetheless, using $u>0$ offers practical benefits: it ensures coverage across different regions of the data space before density-aware sampling begins, and it allows us to accumulate sufficient points to improve the reliability of density estimation needed for effective weight computation. The choice of $u$ is critical to DA-FPS effectiveness. For instance, if set too small, sparse regions may be underrepresented in the training set. This imbalance can result in low model performances in sparse regions of the dataset with a consequential increase in the average error. We speculate that setting $u>0$ may lead to improved approximation guarantees for DA-FPS with respect to minimizing the true weighted fill distance in (2). This is because the initial uniform sampling phase helps ensure that the selected set $\mathcal{L}_{\mathcal{X}}$ has a more balanced representation of the data space, which can lead to more accurate density estimates and weight computations in subsequent iterations. However, formalizing this intuition and deriving precise approximation bounds for DA-FPS with $u>0$ remains an open problem for future research.